OpenStudy (anonymous):

given sec = (radical 5)/2. find the value of tan

OpenStudy (asnaseer):

you are given:\[\sec(\theta)=\frac{\sqrt{5}}{2}\]this can be illustrated by the following diagram:  so, in this diagram:\[\tan(\theta)=\frac{x}{2}\]which means you just need to find x. you Pythagorus theorem for this.